One number is 4 more than another, and their sum is 60. What is the smaller number? If x=the larger number and y=smaller number, then which of the following systems could be used to solve the problem? x+y=60and -x+y=4; x+y=60and x-y=4 or x+y=60and x+y=4
Answer by Alan3354(69443)

One number is 4 more than another, and their sum is 60. What is the smaller number? If x=the larger number and y=smaller number, then which of the following systems could be used to solve the problem?
x+y=60 and -x+y=4;
x+y=60 and x-y=4 or
x+y=60 and x+y=4
----------------
Try each of the 3 pairs, see which one works.